27 +(3x27) divided by 6=
Answer by checkley77(12844)

27 +(3x27) divided by 6=
************************
27+(3*27)/6
27+81/6
27+13.5=13.3 ans.
OR
(27+3*27)/6
(27+81)/6
108/6=18 ans.